Staff: Jacquie, Sam
Weather: Wonderful! Cool and cloudy while at the site. Windy. Clouds started clearing late in the afternoon.

Looking back at the array and trailer tower. Taken ~10am
Notes:
- The towers survived hurricane Hilary!
- We visited the site in the AM before Sam departed (safe travels!)
- Josh arrives tonight.
- We addressed a few priority issues:
tt
- Replaced the batteries and power cycled the DSM's. The beacon is now powered and lit.
- However, the GPS in ttb is still not operating. We replugged, let the GPS sit outside the box for a while, but non of those worked.

t23 DSM
- The cooler was filled with a few inches of rain and the victron was waterlogged. We replaced the victron and everything powered-up.
- I check one of the other coolers just to see if it also had water and found a black window spider and it's babies (not doing that again!).
- Just the CSATS were not coming online so we unplugged/replugged the cables on the CSAT ports.
- We also power cycled the soil mote which was also offline.

t1 DSM
- The CSAT3 and CSAT3A was offline so we unplugged/replugged the cables on their ports. Having the print out in the DSM boxes indicating which port the sensors are on was great!